# Generalization capabilities and robustness of hybrid machine learning models grounded in flow physics compared to purely deep learning models

### v2

@@ IV Discussion
However, as shown in Figs. 5, 7, 10, 15 and 17 the POD-DL model generates the most accurate predictions for either the laminar and turbulent flow. This suggests that, although the unaccounted variations in the $U$ matrix might introduce some errors, these errors are not significant enough to overshadow the advantages of the POD-DL model. This ability to capture the main dynamics of the flow, makes it the most effective model among the models tested, despite the potential inaccuracies introduced by assuming a constant $\bm{U}$ matrix.
@@
Moreover, the ability of POD-DL to handle different datasets with minimal modifications to the forecasting model enhances its versatility. This advantage is particularly valuable in practical applications where datasets vary in resolution. By maintaining a consistent and simplified approach to forecasting, the POD-DL model ensures robustness and adaptability across diverse datasets.
@@ V Conclusions
This combination with the POD, also allows the model to be informed about the physics of the problem, as POD decomposes the spatio-temporal flow dynamics in different POD modes with physical meaning. We believe that this work has demonstrated the potential of combining modal decomposition techniques with forecasting models, to predict the temporal evolution of flow dynamics. Although further research is still needed, especially in the area of probabilistic forecasting, these kind of hybrid models may have a wide variety of applications in both industry and academia.
@@ Acknowledgements.
The authors acknowledge the grants PID2023-147790OB-I00, TED2021- 129774B-C21 and PLEC2022-009235 funded by MCIN/AEI/ 10.13039/501100011033 and by the European Union “NextGenerationEU”/PRTR. The MODELAIR and ENCODING projects have received funding from the European Union’s Horizon Europe research and innovation programme under the Marie Sklodowska-Curie grant agreement No. 101072559 and 101072779, respectively. The results of this publication reflect only the author(s) view and do not necessarily reflect those of the European Union. The European Union can not be held responsible for them.
@@
The code used in this work is available in https://github.com/RAbadiaH/Generalization-capabilities-and-robustness-of-hybrid-machine-learning-models.git. The data of the turbulent flow that was obtained by Mendez et al. Mendez et al. (2020) and supports the findings of this study is openly available in https://github.com/mendezVKI/MODULO/tree/master/download_all_data_exercises.
@@ A.1 Data augmentation
As shown in 14 and 15, when the dataset is three-dimensional, the elements of the $z$ dimension in $\bm{V}$ can be considered independently. Allowing to distinguish $J_{4}$ two-dimensional datasets (16 - 19), each one representing a different time-series, instead of a single three-dimensional dataset.
@@ A.2 Rolling window
All these windows are scaled using the min-max scaling, with the global minimum and maximum values of the original time series. This preprocessing step ensures that the data is normalized to the range $[0,1]$, which helps to improve the training stability and performance of machine learning models, by preventing gradients from diverging or getting stuck.
Appendix B Models architecture and residual autoencoder latent space

### v4

@@ IV Discussion
However, as shown in Fig. 5, Fig. 7, Fig. 10, Fig. 15 and Fig. 17 the POD-DL model generates the most accurate predictions for either the laminar and turbulent flow. This suggests that, although the unaccounted variations in the $U$ matrix might introduce some errors, these errors are not significant enough to overshadow the advantages of the POD-DL model. This ability to capture the main dynamics of the flow, makes it the most effective model among the models tested, despite the potential inaccuracies introduced by assuming a constant $\bm{U}$ matrix.
The poor performance of the models in Fig. 15, at the spatial coordinate $P_{2}$, may be attributed to the limitations of the hidden layer, which consists of either an LSTM or a ConvLSTM, depending on the model. These architectures may struggle to capture the complex dynamics associated with small-scale flow structures. A more effective approach could involve approximating the statistical behavior of these scales. In this context, temporal diffusion models, known for their capability to approximate complex distributions, could serve as a viable alternative to LSTM or ConvLSTM architectures.
@@
Moreover, the ability of POD-DL to handle different datasets with minimal modifications to the forecasting model enhances its versatility. This advantage is particularly valuable in practical applications where datasets vary in resolution. By maintaining a consistent and simplified approach to forecasting, the POD-DL model ensures robustness and adaptability across diverse datasets, as shown in Appendix C where a pre-trained POD-DL is tasked to predict a completely new flow dynamics, with and without retraining the model with the new data.
However, the performance of these data-driven models is intrinsically dependent on the quality and quantity of the training data. Additionally, their predictive accuracy declines as the complexity of the dynamics increases, particularly in turbulent flow regimes. Among the evaluated models, the POD-DL approach exhibits superior performance in managing turbulence compared to the residual autoencoder and VAE. Nevertheless, its effectiveness is highly sensitive to the number of POD modes retained during truncation, where capturing the dynamics of the less energetic modes is particularly challenging due to their chaotic and noisy behavior, often necessitating more advanced models than LSTMs, which in turn may require larger training datasets. This balance between the size of the training dataset and prediction accuracy remains an active area of research.
@@ V Conclusions
This combination with the POD, also allows the model to be informed about the physics of the problem, as POD decomposes the spatio-temporal flow dynamics in different POD modes with physical meaning. We believe that this work has demonstrated the potential of combining modal decomposition techniques with deep learning forecasting models, to predict the temporal evolution of flow dynamics. Further research could explore the integration of more advanced deep learning architectures within the POD-DL methodology, substituting the LSTM layer by a temporal diffusion model Lin et al. (2024), which has shown promising results in weather forecasting Kochkov et al. (2024), as well as the application of these models to a wider range of fluid dynamics problems. Especially those where a larger truncation of POD modes is required to capture the dynamics of interest. Additionally, investigating the use of transfer learning to leverage pre-trained models, from other dynamics, for new flow conditions could be beneficial to a wide variety of applications in both industry and academia.
@@ Acknowledgements.
The authors acknowledge the grants TED2021-129774B-C21 and PLEC2022-009235 funded by MCIN/AEI/10.13039/501100011033 and by the European Union “NextGenerationEU”/PRTR and the grant PID2023-147790OB-I00 funded by MCIU/AEI/10.13039/501100011033/FEDER, UE. The MODELAIR and ENCODING projects have received funding from the European Union’s Horizon Europe research and innovation programme under the Marie Sklodowska-Curie grant agreement No. 101072559 and 101072779, respectively. The results of this publication reflect only the author(s) view and do not necessarily reflect those of the European Union. The European Union can not be held responsible for them.
@@
The code used in this work is available in https://github.com/RAbadiaH/hybrid_and_purely_dl_models.git. The data of the turbulent flow that was obtained by Mendez et al. Mendez et al. (2020) and supports the findings of this study is openly available in https://github.com/mendezVKI/MODULO/tree/master/download_all_data_exercises.
@@ A.1 Data augmentation
As shown in 15 and 16, when the dataset is three-dimensional, the elements of the $z$ dimension in $\bm{V}$ can be considered independently. Allowing to distinguish $J_{4}$ two-dimensional datasets (17 - 20), each one representing a different time-series, instead of a single three-dimensional dataset.
@@ A.2 Rolling window
All these windows are scaled using the min-max scaling, with the global minimum and maximum values of the original time series. This preprocessing step ensures that the data is normalized to the range $[0,1]$, which helps to improve the training stability and performance of machine learning models, by preventing gradients from diverging or getting stuck. This vanishing or exploding gradient problem is also handle with the architectural choice of the models, for example the residual autoencoder utilizes residual connections explicitly designed to mitigate this problem, and the Variational Autoencoder and POD-DL models, make use of a LSTM layer that was incorporated to maintain long-term dependencies, thereby addressing the vanishing gradient issue inherent in recurrent networks. Additionally, a careful weight initialization were employed across all models to further reduce the risk of gradient instability.
Appendix B Stability of POD-DL at different numbers of training samples
This appendix examines the training stability of the POD-DL model when trained with a smaller training set. The dataset chosen for this analysis is the turbulent flow past a cylinder Mendez et al. (2020), due to its inherent complexity.
For this aim, the POD-DL model was independently trained on three distinct training sets, each comprising a different number of samples: $500$, $1000$ and $2800$, where the last one matches the number of samples used to train the residual autoencoder and VAE models, results from this last case can be found in Section III.2.
Figure 21 illustrates the prediction error of the POD-DL model over the same $200$ time-steps analyzed in Section III.2, evaluated across different training sets. The results highlight the model’s training stability, even when fewer training samples are used. In comparison, while the residual autoencoder requires $2800$ samples for training, the POD-DL model maintains relatively consistent performance for smaller training sets, demonstrating robustness to variations in training set size. This robustness arises from the dimensionality reduction achieved by POD, which enables the use of lower-dimensional vectors for training, unlike the high-dimensional snapshots required by the residual autoencoder and VAE models.
Figure 21: Prediction error of the POD-DL model trained with different number of samples (indicated in brackets) for the streamwise velocity component of the turbulent flow past a cylinder. Errors are quantified using (a) the RRMSE and (b) the SSIM defined in (7) and (8), respectively.
Appendix C Results from a pre-trained model
This appendix shows the results from the POD-DL model when is pre-trained on the turbulent flow and task it to forecast the dynamics from a completely different flow. This is done by performing transfer learning, where the weights from a previous training are loaded before forecasting.
Two cases are study in this appendix: (1) forecasting the new dynamics without retraining with snapshots from the new dataset and (2) retraining the POD-DL model with the first $800$ snapshots from the new dataset. This is done to test the generalization capabilities of the POD-DL model to unseen dynamics.
Table 5: Comparison of the original flow from the numerical simulation of the Jet LES and the reconstruction using SVD, retaining the $20$ most energetic modes to reconstruct the flow field.
Original SVD reconstruction
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
Figure 22: Error committed by the POD-DL model with and without retraining, on the streamwise velocity component of the Jet LES. Errors are quantified using (a) the RRMSE and (b) the SSIM defined in (7) and (8), respectively.
Figure 23: Comparison of the velocity histograms between the ground truth and the temporal predictions of the Jet LES obtained from (a) the POD-DL without retraining and (b) the POD-DL with retraining, respectively.
This task was infeasible for the residual autoencoder and the VAE due to their architectures, which is rigidly tied to the dimensions of the snapshots used during the training. Adapting these architectures for snapshots of different dimensions would necessitate significant modifications to the model structure, making it impractical or impossible to reuse the pre-trained weights. This limitation could be addressed using techniques such as Spatial Pyramid Pooling He et al. (2014), which standardizes incoming snapshots to a fixed size.
However, the inherent design of the POD-DL model circumvents this issue entirely, where the dimensionality of the snapshots is irrelevant, as the requirement is solely to truncate to the same number of POD modes as during training.
The dataset chosen to carry out these tests comes from a numerical simulation of an isothermal subsonic jet Brès et al. (2018) originating from a round nozzle with an exit diameter of $D=50$mm. The simulation was performed using the large eddy simulation (LES) flow solver, CharLES, which solves the spatially-filtered compressible Navier-Stokes equations on unstructured grids. The solver employs a finite-volume method combined with third-order Runge-Kutta time integration. The simulation utilized approximately sixteen million control volumes and was conducted for a duration of 2000 acoustic time units ($tc_{\infty}/D$) and at Re $\approx 10^{6}$.
This dataset comprises $5000$ snapshots of the jet’s streamwise velocity, sampled at intervals of $\Delta t=0.02$ acoustic time units. The data is captured on a structured cylindrical output grid that closely approximates the resolution of the underlying LES. The grid extends over a streamwise distance of $30D$ and a radial distance of $6D$, effectively capturing the jet’s spatial dynamics.
To utilize the pre-trained POD-DL model for the turbulent flow, the same number of POD modes must be truncated, specifically $20$. Table 5 presents a comparison between the ground truth snapshots obtained from the numerical simulation of the Jet LES and the SVD reconstruction, retaining only the $20$ most energetic modes.
As illustrated in Fig. 22, there is no significant difference in the results between the cases with and without retraining, suggesting that the prediction accuracy is not substantially impacted by retraining the model using snapshots from the Jet LES dataset. We attribute this to the similarity found in the dynamics of these problems. Although their nature is different, the main dynamics is driven by a periodic solution, where the neural network model has been trained to forecast the flow evolution following this type of dynamics. Further studies should be carried out to test the POD-DL model in quasi-periodic systems, but this remains an open topic for future research.
(a) Ground truth snapshot and predictions at time instant $850$.
(b) Ground truth snapshot and predictions at time instant $900$.
(c) Ground truth snapshot and predictions at time instant $950$.
Figure 24: Snapshots at some representative time instants showing the ground truth streamwise velocity (left column) and the predictions, from the POD-DL model, for the Jet LES. No retrain (middle column) makes reference to the predictions obtain from case (1) and Retrain (right column) from case (2), described at the beginning of Appendix C.
Figure 23 compares the probability distributions of the ground truth dataset with those predicted by the POD-DL model, highlighting the model’s relative accuracy. Notably, the predictions are slightly more accurate when no retraining is applied. This trend is also qualitatively observed in Fig. 24, which provides a visual comparison between the ground truth and predicted snapshots.
Figure 25 presents the loss evolution of the POD-DL model during training across all dataset cases examined in this work: laminar flow, turbulent flow, and Jet LES with and without retraining. Notably, training on the turbulent flow dataset stagnates at a higher loss value compared to other cases, underscoring the increased complexity of this scenario. Additionally, for the Jet LES case, the training dynamics of the POD-DL model appear largely unaffected by retraining.
Figure 25: Loss evolution of the POD-DL model during training for different datasets: laminar flow past a cylinder (Laminar), turbulent flow past a cylinder (Turbulent), Jet LES with retraining (Jet - retrain), and Jet LES without retraining (Jet - no retrain).
Appendix D Models architecture and residual autoencoder latent space
This appendix show the tables describing the detailed architecture for the models presented in Section II: POD-DL (Tab. 6), residual autoencoder (Tabs. 7, 8, 9) and variational autoencoder (Tabs. 10, 11). Note for the autoencoders the architecture described is conditioned on the incoming dataset, which in this case is the laminar flow Le Clainche, Pérez, and Vega (2018), since the architecture of these models directly depends on the resolution of the incoming snapshots.
Figure 26 presents the latent variables generated by the encoder of the residual autoencoder model when applied to the laminar flow. The visualization highlights distinct spatial structures within the flow dynamics, suggesting that the encoder effectively decomposes the flow into its constituent dynamic components.